OpenStudy (anonymous):

simplify (3x+8)(2x+6)

OpenStudy (texaschic101):

Use the FOIL method. (first, outside, inside, last) FIRST -- take the first terms out of each set and multiply them 3x * 2x = 6x^2 OUTSIDE -- take the outside terms in each set and multiply 3x * 6 = 18x INSIDE -- take the inside terms in each set and multiply 8 * 2x = 16x LAST -- take the last numbers out of each set and multiply 8 * 6 = 48 now put them all together.. 6x^2 + 18x + 16x + 48 (combine like terms) 6x^2 + 34x + 48 I don't know if you need to, but this can be factored 2(3x^2 + 17x + 24)